OpenStudy (anonymous):

a=1\2bh resolve for b

OpenStudy (australopithecus):

is that, \[a = \frac{1}{2bh}\]

rvc (rvc):

cross multiply

OpenStudy (australopithecus):

multiply both sides by h then divide both sides by a

rvc (rvc):

2bh= 1/a b=1/2ah

OpenStudy (australopithecus):

or conversely multiply both sides by (1/a) for the second step it is the same thing

OpenStudy (anonymous):

thaks

rvc (rvc):

hope u got the method! @tito23

OpenStudy (anonymous):

is this the answer 2a=bh. b=2a\h

rvc (rvc):